KrazyKat02:

Nina's monthly budget is $2,250. Every month she makes the following payments: $175 for insurance $129 for utilities $283 for taxes Which is the highest house payment she can afford? A. House A - monthly payment: $1,580.89 B. House B - monthly payment: $1,790.25 C. House C - monthly payment: $1,989.23

LadyRavyRay:

2250-(all the utilities)=2250-(283+129+175)=2250-(587) Can you calculate the rest?

KrazyKat02:

Yeah

KrazyKat02:

I got $1,663. But that isn't one of my answer choices.

LadyRavyRay:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @KrazyKat02 I got $1,663. But that isn't one of my answer choices. \(\color{#0cbb34}{\text{End of Quote}}\) You're close to the answer. The total costs that Nina would have to pay is the sum of the cost of insurance, utilities, and taxes. C = $175 + $129 + $283 = $587 The highest house payment she could afford is then the difference between her budget and the costs, P = $2,250 - $587 = $1663 So, what's the answer?

xxDeppressionxx:

The answer would be (B.) House B because all the rest of the monthly payment are way over her budget. C = $175 + $129 + $283 = $587 P = $2,250 - $587 = $1663

KrazyKat02:

Um, I feel like I should know it, but I'm having a brain fart moment right now.

KrazyKat02:

Ooooooh

KrazyKat02:

Okay. Thank you guys so much!

xxDeppressionxx:

Np

LadyRavyRay:

Yeah, the answer is B. It's not going over the budget.